9.50*10^22 formula units of lead (II) nitrate is how many miles?
1
Expert's answer
2020-04-15T01:51:24-0400

Answer on Question #109610, Chemistry, General Chemistry

9.50*10^22 formula units of lead (II) nitrate is how many moles?


Solution:

Number of moles of lead (II) nitrate can be determine as:

n (Pb(NO3)2) = N/Na, where N = 9.50*1022; Na is Avogadro's number which equals 6.022*1023 mol-1.

Thus:

n (Pb(NO3)2) = 9.50*1022/6.022*1023 = 0.158 (moles)


Answer: n (Pb(NO3)2) = 0.158 moles.